Solve the equation. log, x + log, (x - 7) = 3 O {5} O {-1,8} O {8} O {1} O {-8, 1)
Answer #1

02 3 = (-) ها +18 ها 3 = [2? - g [ x Cx) < 3 ت 82) د [72-7x] = .2 2 * - .12 2 -77 - 8 = 0 د ن = (8 – 8) را + ) 8 راس = X